# Semi-Supervised Deep Sobolev Regression: Estimation, Variable Selection and Beyond

### v1

@@ 1.3 Preliminaries
Similarly, one can define the $L^{2}(\nu_{X})$ inner-product and norm, by
[…]
Furthermore, define the density ratio between $\nu_{X}$ and $\mu_{X}$ by $r(x)=q(x)/p(x)$. Suppose the density ratio is uniformly upper- and lower-bounded, that is,
Then it is straightforward to verify that
@@
Standard $H^{1}(\nu_{X})$ Sobolev semi-inner-product and semi-norm are given, respectively, by
$(u,v)_{H^{1}(\nu_{X})}=\int_{\Omega}\sum_{k=1}^{d}D_{k}uD_{k}vd\nu_{X}\quad\text{and}\quad|u|_{H^{1}(\nu_{X})}^{2}=(u,u)_{H^{1}(\nu_{X})}.$
For two functions $u,v\in H^{1}(\nu_{X})$, we define the inner products between their gradients $\nabla u=(D_{k}u)_{k=1}^{d}$ and $\nabla v=(D_{k}v)_{k=1}^{d}$ as
$(\nabla u,\nabla v)_{L^{2}(\nu_{X})}=\sum_{k=1}^{d}(u_{k},v_{k})_{H^{1}(\nu_{X})}=(u,v)_{H^{1}(\nu_{X})}.$
Then the corresponding norm is defined as $\|\nabla u\|_{L^{2}(\nu_{X})}=|u|_{H^{1}(\nu_{X})}$.
Definition 1.1 (Continuous functions space).
@@
where $D^{\alpha}=D_{1}^{\alpha_{1}}\cdots D_{d}^{\alpha_{d}}$ with multi-index $\alpha=(\alpha_{1},\ldots,\alpha_{d})^{T}\in{\mathbb{N}}^{d}$.
Definition 1.2 (Deep ReQU neural network).
@@
$\psi(x)=T_{L}(\varrho(T_{L-1}(\cdots\varrho(T_{0}(x))\cdots))),$
where the ReQU activation function $\varrho(x)=(\max\{x,0\})^{2}$ is applied component-wisely and $T_{\ell}(x):=A_{\ell}x+b_{\ell}$ is an affine transformation with $A_{\ell}\in{\mathbb{R}}^{N_{\ell+1}\times N_{\ell}}$ and $b_{\ell}\in{\mathbb{R}}^{N_{\ell}}$ for $\ell=0,\ldots,L$. In this paper, we consider the case $N_{0}=d$ and $N_{L+1}=1$. The number $L$ is called the depth of neural networks. Additionally, $S:=\sum_{\ell=0}^{L}(\|A_{\ell}\|_{0}+\|b_{\ell}\|_{0})$ represents the total number of non-zero weights within the neural network. We denote by ${\mathcal{N}}(L,S)$ the set of functions implemented by ReQU-neural networks with depth at most $L$ and the number of non-zero weights at most $S$.
Definition 1.3 (VC-dimension).
A set $\{x_{i}\}_{i=1}^{m}$ is said to be shattered by ${\mathcal{F}}$ when |$|\{(f(x_{1}),\ldots,f(x_{m})):f\in{\mathcal{F}}\}|=2^{m}$. The Vapnik-Chervonenkis dimension of ${\mathcal{F}}$, denoted $\operatorname{VCdim}({\mathcal{F}})$, is the size of the largest set that can be shattered by ${\mathcal{F}}$, that is, $\operatorname{VCdim}({\mathcal{F}})=\max\{m:\Pi_{{\mathcal{F}}}(m)=2^{m}\}$. For a class ${\mathcal{F}}$ of real-valued functions, we define $\operatorname{VCdim}({\mathcal{F}})=\operatorname{VCdim}(\operatorname{sign}({\mathcal{F}}))$.
For simplicity of notation, for a hypothesis class ${\mathcal{F}}\subseteq H^{1}(\nu_{X})$, we define the class of the first-order partial derivatives as $D_{k}{\mathcal{F}}=\{D_{k}f:f\in{\mathcal{F}}\}$ for $1\leq k\leq d$. We also define their VC-dimension as $\operatorname{VCdim}(D_{k}{\mathcal{F}})$. Further, denote by $\operatorname{VCdim}(\nabla{\mathcal{F}})=\max_{1\leq k\leq d}\operatorname{VCdim}(D_{k}{\mathcal{F}})$.
1.4 Notation
@@
1.5 Organization
The remainder of the article is organized as follows. We commence with a review of related work in Section 2. Subsequently, we outline the deep Sobolev penalized regression and propose the semi-supervised estimator in Section 3. We present the convergence rate analysis for the regression in Section 4 and for the derivative estimation in Section 5. In Section 6, we apply our method to nonparametric variable selection and inverse source problems, and provide an abundance of numerical studies. The article concludes with a few summarizing remarks in Section 7. All technical proofs are relegated to the supplementary material.
2 Related Work
@@
2.1 Nonparametric Derivative Estimation
As previously indicated, the necessity to estimate derivatives arises in various application contexts. Among the simplest and most forthright methods for derivative estimation is the direct measurement of derivatives. For example, in the field of economics, estimating cost functions (Shephard, 1981) frequently involves data on a function and its corresponding set of derivatives. A substantial volume of literature (Florens, Ivaldi and Larribeau, 1996; Peter Hall and Adonis Yatchew, 2007; Hall and Yatchew, 2010) considers this scenario by reverting to a corresponding regression model:
@@ 2.1 Nonparametric Derivative Estimation
To estimate derivatives with noisy measurements only on function values, researchers have put forward nonparametric derivative estimators (Newell and Einbeck, 2007). Nonparametric derivative estimation encompasses four primary approaches: local polynomial regression (De Brabanter et al., 2013), smoothing splines (Heckman and Ramsay, 2000), kernel ridge regression (Liu and Li, 2023), and difference quotients (Müller, Stadtmüller and Schmitt, 1987; Liu and De Brabanter, 2018; Liu and Brabanter, 2020). Among these approaches, the first three are categorized as plug-in derivative estimators. In this article, we present a review of these plug-in approaches using the one-dimensional case as an illustrative example.
2.1.1 Local Polynomial Regression
In standard polynomial regression, a single polynomial function is used to fit the data. One of the main challenges with this method is the need to use high-order polynomials to achieve a more accurate approximation. However, high-order polynomials may be oscillative in some regions, which is known as Runge phenomenon (Epperson, 1987). To repair the drawbacks of the polynomial regression, a natural way is to employ the low-degree polynomial regression locally, which is called local polynomial regression (Fan and Gijbels, 1996). Derivative estimation using local polynomial regression was first proposed by De Brabanter et al. (2013). Let $K$ be a kernel function and $h$ be the bandwidth controlling the smoothness. We assign a weight $K((X_{i}-x)/h)$ to the point $(X_{i},Y_{i})$, leading to the following weighted least squares problem:
@@ 2.1.1 Local Polynomial Regression
Herr the kernel $K$ should decay fast enough to eliminate the impact of a remote data point. Denote by $\{\widehat{\beta}_{\ell}\}_{\ell=0}^{p}$ the estimator obtained by (4). The estimated regression curve at point $x$ is given by $\widehat{f}(x)=\sum_{\ell=0}^{p}\widehat{\beta}_{\ell}(x)(X_{i}-x)^{\ell}$. Further, according to Taylor’s theorem, the estimator of the first order derivative $f_{0}^{\prime}$ at point $x$ is given by $\widehat{f}^{\prime}(x)=\widehat{\beta}_{1}(x)$. Masry (1996a, b) established the uniform strong consistency and the convergence rates for the regression function and its partial derivatives. Derivative estimation using local polynomial regression in multivariate data has been discussed in Amiri and Thiam (2018).
2.1.2 Smoothing Splines
Extensive research has been conducted on the use of smoothing splines in nonparametric regression (Wahba, 1990; Green and Silverman, 1993; Eubank, 1999; Wasserman, 2006). This method starts from the minimization of a penalized least squares risk
@@ 2.1.2 Smoothing Splines
where the first term encourages the fitting of estimator to data, the second term penalizes the roughness of the estimator, and the smoothing parameter $\gamma>0$ controls the trade-off between the two conflicting goals. The minimizer $\widehat{f}$ of (5) is an estimator of the regression function $f_{0}$, which is called cubic smoothing spline. The plug-in derivative estimator $\widehat{f}^{\prime}$ is a direct estimate of the derivative $f_{0}^{\prime}$ of the regression function. This idea has been pursued by Stone (1985); Heckman and Ramsay (2000). In the perspective of theoretical analysis, Stone (1985) shows that spline derivative estimators can achieve the optimal rate of convergence, and Zhou and Wolfe (2000) studies local asymptotic properties of derivative estimators.
2.1.3 Kernel Ridge Regression
Kernel ridge regression is a technique extensively employed in the domain of nonparametric regression. Liu and Li (2023) introduced a plug-in kernel ridge regression estimator for derivatives of the regression function, establishing a nearly minimax convergence rate for univariate function classes within a random-design setting. Further expanding upon this method, Liu, Li and Li (2023) applied it to multivariate regressions under the smoothing spline ANOVA model and established minimax optimal rates. Additionally, Liu, Li and Li (2023) put forth a hypothesis testing procedure intended to determine whether a derivative is zero.
2.2 Nonparametric Regression using Deep Neural Network
In comparison to the nonparametric methods mentioned above, deep neural networks (Goodfellow, Bengio and Courville, 2016) also stand out as a formidable technique employed within machine learning and nonparametric statistics. Rigorous of the convergence rate analysis have been established for deep nonparametric regression (Bauer and Kohler, 2019; Nakada and Imaizumi, 2020; Schmidt-Hieber, 2020; Kohler and Langer, 2021; Farrell, Liang and Misra, 2021; Kohler, Krzyżak and Langer, 2022; Jiao et al., 2023), but derivative estimation using deep neural networks remained an open problem prior to this paper, even though the derivative of the regression estimate is of great importance as well.
Unfortunately, estimating derivatives is not always a by-product of function estimation. Indeed, the basic mathematical analysis (Tao, 2022, Section 3.7) shows that, even if estimators $\{f_{n}\}_{n\geq 1}$ converge to the regression function $f_{0}$, the convergence of plug-in derivative estimators $\{\nabla f_{n}\}_{n\geq 1}$ is typically not guaranteed. To give a counterexample, we consider the functions $f_{n}:I\rightarrow{\mathbb{R}},x\mapsto n^{-1}\sin(nx)$, and let $f_{0}:I\rightarrow{\mathbb{R}}$ be the zero function $f_{0}(x)=0$, where $I:=[0,2\pi]$. Then $\|f_{n}-f_{0}\|_{L^{p}(I)}\rightarrow 0$ as $n\rightarrow 0$, but $\lim_{n\rightarrow\infty}\|f_{n}^{\prime}-f_{0}^{\prime}\|_{L^{p}(I)}\neq 0$ for each $1\leq p\leq\infty$.
@@ 2.2 Nonparametric Regression using Deep Neural Network
2.3 Nonparametric Vairable Selection
@@
Data collected in real-world applications tend to be high-dimensional, although only a subset of the variables within the covariate vector may genuinely exert influence. Consequently, variable selection becomes critical in statistics and machine learning as it both mitigates computational complexity and enhances the interpretability of the model. However, traditional methods for variable selection have been primarily focused on linear or additive models and do not readily extend to nonlinear problems. One inclusive measure of the importance of each variable in a nonlinear model is its corresponding partial derivatives. Building on this concept, a series of works (Rosasco et al., 2010; Mosci et al., 2012; Rosasco et al., 2013) introduced sparse regularization to kernel ridge regression for variable selection. They have consequently devised a feasible computational learning scheme and developed consistency properties of the estimator. However, the theoretical analysis is limited to reproducing kernel Hilbert space and cannot be generalized to deep neural network-based methods.
2.4 Semi-Supervised Learning
@@
Semi-supervised learning has recently gained significant attention in statistics and machine learning (Zhu and Goldberg, 2009; Van Engelen and Hoos, 2020). The basic setting of semi-supervised learning is common in many practical applications where the label is often more difficult or costly to collect than the covariate vector. Therefore, the fundamental question is how to design appropriate learning algorithms to fully exploit the value of unlabeled data. In the past years, significant effort has been devoted to studying the algorithms and theory of semi-supervised learning (Zhang, 2000; Belkin, Niyogi and Sindhwani, 2006; Wasserman and Lafferty, 2007; Azriel et al., 2022; Livne, Azriel and Goldberg, 2022; Song, Lin and Zhou, 2023; Deng et al., 2023). The most related work is Belkin, Niyogi and Sindhwani (2006), whose main idea is to introduce an unlabeled-data-driven regularization term to the loss function. Specifically, Belkin, Niyogi and Sindhwani (2006) employ a manifold regularization to incorporate additional information about the geometric structure of the marginal distribution, where the regularization term is estimated on the basis of unlabeled data. In addition, our method does not require the distribution of the unlabeled data to be aligned with the marginal distribution of the labeled data exactly, which expands the applicability scenarios.
2.5 Ill-Posed Problems and Tikhonov Regularization
Sobolev penalized least-squares regression exhibits close ties to inverse problems. In the context of estimating derivatives based on noisy measurements using difference quotients, the problem often becomes ill-posed, meaning that even minor amounts of noise can result in significant estimation errors. Sobolev penalized least-squares regression offers a reliable approach to tackle this ill-posed problem. For example, Hu, Shu and Zou (2018) applied Sobolev penalty least-squares to the inverse source problem, and this strategy can also been applied in Jin et al. (2023) to estimate the gradient of solution. Sobolev penalty serves as a crucial regularization strategy for inverse problems, as evidenced by Tikhonov regularization, which formulates an ill-posed problem as a Sobolev penalized least-squares regression with PDE constraints (Nickl, van de Geer and Wang, 2020).
3 Deep Sobolev Regression
@@
In this section, we present an in-depth examination of Sobolev penalized least-squares regression as implemented through deep neural networks. Initially, we incorporate the $H^{1}$-semi-norm penalty into the least-squares risk. Subsequently, we delineate the deep Sobolev regressor as referenced in Section 3.1, followed by an introduction to the semi-supervised Sobolev regressor elaborated in Section 3.2.
@@
Here the regularization parameter $\lambda$ governs the delicate equilibrium between conflicting objectives: data fitting and smoothness. Specifically, when $\lambda$ nearly or entirely vanishes, (6) aligns with the standard population least-squares risk. Conversely, as $\lambda$ approaches infinity, the minimizer of (6) tends towards a constant estimator. For the joint distribution $\mu$ of $(X,Y)$, let $\mu_{X}$ denote the margin distribution of $X$. According to (1), one obtains easily
@@
[…]
Sobolev penalized regression can be interpreted as a PDE-based smoother of the regression function $f_{0}$. Let $f^{\lambda}$ denote a solution to the quadratic optimization problem (6). Some standard calculus of variations (Brenner and Scott, 2008; Evans, 2010) show that, if the minimizer $f^{\lambda}$ has square integrable second derivatives, then $f^{\lambda}$ solves the following second-order linear elliptic equation with homogeneous Neumann boundary condition:
$-\lambda\Delta f^{\lambda}+f^{\lambda}&=f_{0},&&\text{in}~{}\Omega,\\ \nabla f^{\lambda}\cdot\bm{n}&=0,&&\text{on}~{}\partial\Omega,\end{aligned}\right.$
In the context of partial differential equations (PDE), the variational problem (6) is called Ritz method (Brenner and Scott, 2008, Remark 2.5.11). The following lemma shows the uniqueness of solution to the above equation.
Lemma 3.1 (Existence and uniqueness of population risk minimizer).
@@
In practical application, the data distribution $\mu$ in (6) remains unknown, making the minimization of population risk (6) unattainable. The goal of regression is to estimate the function $f_{0}$ from a finite set of data pairs ${\mathcal{D}}=\{(X_{i},Y_{i})\}_{i=1}^{n}$ which are independently and identically drawn from $\mu$, that is,
@@
3.1 Deep Sobolev regressor

### v2

@@ I-C Preliminaries and notations
$\|u\|_{L^{2}(\mu_{X})}^{2}=(u,u)_{L^{2}(\mu_{X})}.$
Similarly, one can define the $L^{2}(\nu_{X})$ inner-product and norm. Furthermore, define the density ratio between $\nu_{X}$ and $\mu_{X}$ by $r(x)=q(x)/p(x)$. Suppose the density ratio is uniformly upper- and lower-bounded, that is, $\kappa:=\sup_{x\in\Omega}|r(x)|<\infty$ and $\zeta:=\inf_{x\in\Omega}|r(x)|>0$. Then it is straightforward to verify that
@@
For two functions $u,v\in H^{1}(\nu_{X})$, the inner products between their gradients is defined as
$(\nabla u,\nabla v)_{L^{2}(\nu_{X})}=\int_{\Omega}\sum_{k=1}^{d}D_{k}uD_{k}vd\nu_{X}.$
Definition 1 (Continuous functions space).
@@
where $D^{\alpha}=D_{1}^{\alpha_{1}}\cdots D_{d}^{\alpha_{d}}$ with $\alpha=(\alpha_{1},\ldots,\alpha_{d})^{T}\in{\mathbb{N}}^{d}$.
Next, we introduce the concept of a deep neural network. While deep ReLU neural networks have shown empirical success in nonparametric regression tasks, they are not suitable for scenarios where derivatives of the network are required in the objective function [46]. This limitation arises from the piecewise linear nature of the ReLU activation function, which results in a lack of continuous derivatives. In contrast, the Rectified Quadratic Unit (ReQU) activation function, defined as the square of the ReLU function, possesses a continuous first derivative. This characteristic allows us to incorporate the deep ReQU neural network in the SDORE framework, thereby expanding the possibilities for the simultaneous estimation of regression values and their derivatives.
Definition 2 (Deep ReQU neural network).
@@
$\psi(x)=T_{L}(\varrho(T_{L-1}(\cdots\varrho(T_{0}(x))\cdots))),$ (6)
where the ReQU activation function $\varrho(x)=(\max\{x,0\})^{2}$ is applied component-wisely and $T_{\ell}(x):=A_{\ell}x+b_{\ell}$ is an affine transformation with $A_{\ell}\in{\mathbb{R}}^{N_{\ell+1}\times N_{\ell}}$ and $b_{\ell}\in{\mathbb{R}}^{N_{\ell}}$ for $\ell=0,\ldots,L$. In this paper, we consider the case $N_{0}=d$ and $N_{L+1}=1$. The number $L$ is called the depth of the neural network, and the number $\max_{1\leq\ell\leq L}N_{\ell}$ is called the width of the neural network.Additionally, $\sum_{\ell=0}^{L}(\|A_{\ell}\|_{0}+\|b_{\ell}\|_{0})$ represents the total number of non-zero weights within the neural network. The space of deep ReQU neural networks with given network architecture is defined as
@@
To measure the complexity of a function class, we next introduce the empirical covering number.
Definition 3 (Empirical covering number).
is called the $L^{p}({\mathcal{D}})$ $\delta$-covering number of ${\mathcal{F}}$.
We now introduce some basic notations. The set of positive integers is denoted by +=${\mathbb{N}}_{+}=\{1,2,\ldots\}$. Denote ={0}+${\mathbb{N}}=\{0\}\cup{\mathbb{N}}_{+}$ for convenience. For a positive integer $m\in{\mathbb{N}}_{+}$, let $[m]$ denote the set $\{1,\ldots,m\}$. We employ the notations $A\lesssim B$ and $B\gtrsim A$ to signify that there exists an absolute constant $c>0$ such that $A\leq cB$.
I-D Organization
The remainder of the article is organized as follows. We commence with a review of related work in Section II. Subsequently, we outline the deep Sobolev penalized regression and propose the semi-supervised estimator in Section III. We present the convergence rate analysis for the regression in Section IV and for the derivative estimation in Section V. In Section VI, we apply our method to nonparametric variable selection, and provide an abundance of numerical studies. The article concludes with a few summarizing remarks in Section VII. All technical proofs are relegated to the supplementary material.
II Related Work
@@
II-A Nonparametric Derivative Estimation
As previously indicated, the necessity to estimate derivatives arises in various application contexts. Among the simplest and most forthright methods for derivative estimation is the direct measurement of derivatives. For example, in the field of economics, estimating cost functions [17] frequently involves data on a function and its corresponding set of derivatives. A substantial volume of literature [47, 48, 49] considers this scenario by reverting to a corresponding regression model:
@@ II-A Nonparametric Derivative Estimation
To estimate derivatives with noisy measurements only on function values, researchers have put forward nonparametric derivative estimators [50]. Nonparametric derivative estimation encompasses four primary approaches: local polynomial regression [27], smoothing splines [28], kernel ridge regression [29], and difference quotients [30, 51, 52]. Among these approaches, the first three are categorized as plug-in derivative estimators. In this article, we present a review of these plug-in approaches using the one-dimensional case as an illustrative example.
II-A1 Local Polynomial Regression
In standard polynomial regression, a single polynomial function is used to fit the data. One of the main challenges with this method is the need to use high-order polynomials to achieve a more accurate approximation. However, high-order polynomials may be oscillative in some regions, which is known as Runge phenomenon [53]. To repair the drawbacks of the polynomial regression, a natural way is to employ the low-degree polynomial regression locally, which is called local polynomial regression [54]. Derivative estimation using local polynomial regression was first proposed by [27]. Let $K$ be a kernel function and $h$ be the bandwidth controlling the smoothness. We assign a weight $K((X_{i}-x)/h)$ to the point $(X_{i},Y_{i})$, leading to the following weighted least-squaress problem:
@@ II-A1 Local Polynomial Regression
Herr the kernel $K$ should decay fast enough to eliminate the impact of a remote data point. Denote by $\{\widehat{\beta}_{\ell}\}_{\ell=0}^{p}$ the estimator obtained by (7). The estimated regression curve at point $x$ is given by $\widehat{f}(x)=\sum_{\ell=0}^{p}\widehat{\beta}_{\ell}(x)(X_{i}-x)^{\ell}$. Further, according to Taylor’s theorem, the estimator of the first order derivative $f_{0}^{\prime}$ at point $x$ is given by $\widehat{f}^{\prime}(x)=\widehat{\beta}_{1}(x)$[55, 56] established the uniform strong consistency and the convergence rates for the regression function and its partial derivatives. Derivative estimation using local polynomial regression in multivariate data has been discussed in [57].
II-A2 Smoothing Splines
Extensive research has been conducted on the use of smoothing splines in nonparametric regression [41, 58, 59, 2]. This method starts from the minimization of a penalized least-squaress risk
@@ II-A2 Smoothing Splines
where the first term encourages the fitting of estimator to data, the second term penalizes the roughness of the estimator, and the smoothing parameter $\gamma>0$ controls the trade-off between the two conflicting goals. The minimizer $\widehat{f}$ of (8) is an estimator of the regression function $f_{0}$, which is called cubic smoothing spline. The plug-in derivative estimator $\widehat{f}^{\prime}$ is a direct estimate of the derivative $f_{0}^{\prime}$ of the regression function. This idea has been pursued by [44, 28]. In the perspective of theoretical analysis, [44] shows that spline derivative estimators can achieve the optimal rate of convergence, and [60] studies local asymptotic properties of derivative estimators.
II-A3 Kernel Ridge Regression
Kernel ridge regression is a technique extensively employed in the domain of nonparametric regression. [29] introduced a plug-in kernel ridge regression estimator for derivatives of the regression function, establishing a nearly minimax convergence rate for univariate function classes within a random-design setting. Further expanding upon this method, [33] applied it to multivariate regressions under the smoothing spline ANOVA model and established minimax optimal rates. Additionally, [33] put forth a hypothesis testing procedure intended to determine whether a derivative is zero.
II-B Nonparametric Regression using Deep Neural Network
In comparison to the nonparametric methods mentioned above, deep neural networks [32] also stand out as a formidable technique employed within machine learning and nonparametric statistics. Rigorous of the convergence rate analysis have been established for deep nonparametric regression [34, 35, 36, 37, 38, 39, 40], but derivative estimation using deep neural networks remained an open problem prior to this paper, even though the derivative of the regression estimate is of great importance as well.
Unfortunately, estimating derivatives is not always a by-product of function estimation. Indeed, the basic mathematical analysis [61, Section 3.7] shows that, even if estimators $\{f_{n}\}_{n\geq 1}$ converge to the regression function $f_{0}$, the convergence of plug-in derivative estimators $\{\nabla f_{n}\}_{n\geq 1}$ is typically not guaranteed. To give a counterexample, we consider the functions $f_{n}:I\rightarrow{\mathbb{R}},x\mapsto n^{-1}\sin(nx)$, and let $f_{0}:I\rightarrow{\mathbb{R}}$ be the zero function $f_{0}(x)=0$, where $I:=[0,2\pi]$. Then $\|f_{n}-f_{0}\|_{L^{p}(I)}\rightarrow 0$ as $n\rightarrow 0$, but $\lim_{n\rightarrow\infty}\|f_{n}^{\prime}-f_{0}^{\prime}\|_{L^{p}(I)}\neq 0$ for each $1\leq p\leq\infty$.
@@ II-B Nonparametric Regression using Deep Neural Network
II-C Nonparametric Vairable Selection
@@
Data collected in real-world applications tend to be high-dimensional, although only a subset of the variables within the covariate vector may genuinely exert influence. Consequently, variable selection becomes critical in statistics and machine learning as it both mitigates computational complexity and enhances the interpretability of the model. However, traditional methods for variable selection have been primarily focused on linear or additive models and do not readily extend to nonlinear problems. One inclusive measure of the importance of each variable in a nonlinear model is its corresponding partial derivatives. Building on this concept, a series of works [23, 24, 25] introduced sparse regularization to kernel ridge regression for variable selection. They have consequently devised a feasible computational learning scheme and developed consistency properties of the estimator. However, the theoretical analysis is limited to reproducing kernel Hilbert space and cannot be generalized to deep neural network-based methods.
II-D Semi-Supervised Learning
@@
Semi-supervised learning has recently gained significant attention in statistics and machine learning [62, 63]. The basic setting of semi-supervised learning is common in many practical applications where the label is often more difficult or costly to collect than the covariate vector. Therefore, the fundamental question is how to design appropriate learning algorithms to fully exploit the value of unlabeled data. In the past years, significant effort has been devoted to studying the algorithms and theory of semi-supervised learning [64, 65, 66, 67, 68, 69, 70]. The most related work is [65], whose main idea is to introduce an unlabeled-data-driven regularization term to the loss function. Specifically, [65] employ a manifold regularization to incorporate additional information about the geometric structure of the marginal distribution, where the regularization term is estimated on the basis of unlabeled data. In addition, our method does not require the distribution of the unlabeled data to be aligned with the marginal distribution of the labeled data exactly, which expands the applicability scenarios.
III Deep Sobolev Regression
@@
In this section, we present an in-depth examination of Sobolev penalized least-squares regression as implemented through deep neural networks. Initially, we incorporate the $H^{1}$-semi-norm penalty into the least-squares risk. Subsequently, we delineate the deep Sobolev regressor as referenced in Section III-A, followed by an introduction to the semi-supervised Sobolev regressor elaborated in Section III-B.
@@
Here the regularization parameter $\lambda>0$ governs the delicate equilibrium between conflicting objectives: data fitting and smoothness. Specifically, when $\lambda$ nearly or entirely vanishes, (9) aligns with the standard population least-squares risk. Conversely, as $\lambda$ approaches infinity, the minimizer of (9) tends towards a constant estimator. For the joint distribution $\mu$ of $(X,Y)$, let $\mu_{X}$ denote the margin distribution of $X$. According to (1), one obtains easily
@@
Sobolev penalized regression can be interpreted as a PDE-based smoother of the regression function $f_{0}$. Let $f^{\lambda}$ denote a solution to the quadratic optimization problem (9). Some standard calculus of variations [71, 72] show that, if the minimizer $f^{\lambda}$ has square integrable second derivatives, then $f^{\lambda}$ solves the following second-order linear elliptic equation with homogeneous Neumann boundary condition:
$-\lambda\Delta f^{\lambda}+f^{\lambda}&=f_{0},&&\text{in}~{}\Omega,\\ \nabla f^{\lambda}\cdot{\mathbf{n}}&=0,&&\text{on}~{}\partial\Omega.\end{aligned}\right.$
In the context of partial differential equations (PDE), the variational problem (9) is called Ritz method [71, Remark 2.5.11]. The following lemma shows the uniqueness of solution to the above PDE.
Lemma III.1 (Existence and uniqueness of population risk minimizer).
@@
In practical applications, the data distribution $\mu$ in (9) remains unknown, making the minimization of population risk (9) unattainable. The goal of regression is to estimate the function $f_{0}$ from a finite set of data pairs ${\mathcal{D}}=\{(X_{i},Y_{i})\}_{i=1}^{n}$ which are independently and identically drawn from $\mu$, that is,
@@
III-A Deep Sobolev regressor